OpenStudy (anonymous):

solve the following using midpoint rule and simpson rule.. integral of (x/(1+x^2)), [0,2], n=10

OpenStudy (mathmale):

Would be happy to help, but invite you to be involved as much as you possibly can be. In this problem we're aiming to find the area between the graph of the given function on the interval that begins with a and ends with b. We'll divide up that area into n vertical strips. Then the width of each vertical strip of area is \[\Delta x=\frac{ b-a }{ n }\] Would you calculate this width, please?

OpenStudy (anonymous):

Sure. delta x= 2-0/10=1/5

OpenStudy (mathmale):

That's correct. Good for you. One suggestion, however: Please enclose that x-2 inside parentheses: (2-0)/10. Next, we need to find n+1 x-values at which to evaluate f(x). The first one will be a = 0, and the second b = 2. What are the 9 intermediate values?

OpenStudy (anonymous):

Is it starts with 0.1, 0.3...?

OpenStudy (mathmale):

I'm very sorry. I meant that the first x-value would be 0 and the LAST would be 2. If the first x-value is 0, then our set of x-values starts out as 0, 0.2, 0.4, ... , and ends with with ... ??

OpenStudy (mathmale):

Remember, delta x is 0.2.

OpenStudy (anonymous):

set of x is 0,0.2,0.4,...,2 intermediate value is 0.1,0.3,0.5,....1.9 right?

OpenStudy (mathmale):

There are n+1 such x- values, meaning 11 x-values. They are, as you said above, 0,0.2,0.4,...,2. (Save the intermediate values for the "midpoint" approach, which we'll tackle later. For now, we're focusing on the use of Simpson's Rule.

OpenStudy (mathmale):

lala: Have you any references available, e. g., your course textbook, access to Internet searches, or the like? If so, could you please look up the formula for Simpson's Rule, if you don't know it already?

OpenStudy (anonymous):

Formula: h/3(y0+4y1+y2)

OpenStudy (mathmale):

that formula will be our starting point, towards approximating the area under the curve between x=0 and x=2. Your h/3(y0+4y1+y2) is all right as a starting point. Again, please enclose h/3 within parentheses so that there's no ambiguity about h/3 being a fraction. Since n=10, you will have n+1, or 11, x-values and 11 corresponding y values. How do you normally go about evaluating a function such as f(x) = x / (1+x^2) at various x-values?

OpenStudy (mathmale):

Starting with your (h/3(y0+4y1+y2+ ..... ), let's write Simpson's Rule more formally:

OpenStudy (mathmale):

\[A=\frac{ \Delta x}{ 3 }(1*f(x _{0}) + 4*f(x _{1}) + 2*f(x _{2}) + ,,, +4*f(x _{9})+1*f(x _{10}))\]

OpenStudy (mathmale):

I'm assuming you've seen this format before. Right?

OpenStudy (anonymous):

Yes!

OpenStudy (mathmale):

And we already know that delta x is 1/5, or 0.2, and that the first x-value is zero, and so on. So, what is the value of the function at x0 = 0?

OpenStudy (anonymous):

Is it 0?

OpenStudy (mathmale):

Yes. Basically you'll need to create a table with those 11 x-values in the left column and the values of the function x/(1+x^2) in the second column. I asked you before how you 'd go about calculating those function values.

OpenStudy (mathmale):

Lala, I have begun such a table, using Excel on my computer. Are you able to finish this table, that is, are you able to calculate the remaining function values?

OpenStudy (anonymous):

yes!

OpenStudy (mathmale):

OK, then please go back to to the formula I typed in earlier. Area = A = [(delta x)/3]*(1*f(0 + 4*f(0.2) + ................ ).

OpenStudy (mathmale):

could you describe in words what you have to do to finish this problem?

OpenStudy (anonymous):

substitute values of f(x) from x0 to x11 into the formula

OpenStudy (mathmale):

Let me re-word that a bit: "substitute the 11 x-values into the formula f(x), one by one.

OpenStudy (mathmale):

Right. And then what?

OpenStudy (anonymous):

Yes! Then, calculate the answer by multiply (delta x/3) with the total of f(x)

OpenStudy (mathmale):

Hint: remember those coefficients, {1,4,2,4,2,4, ... 4,1}? Not "total of f(x)," but "the sum of the products 1*f(0),4*f(0.2), 2*f(0.4), and so on."

OpenStudy (anonymous):

Yes, that's correct^_^

OpenStudy (mathmale):

Organizing all this info in a table should be helpful. I've come up with the following table and ask you whether or not you feel comfortable completing it:

OpenStudy (mathmale):

Can you view this table? If so, what's the next thing you'd do to complete this table?

OpenStudy (anonymous):

yes. now I have all the coeff*functions value. I am going to total them up

OpenStudy (mathmale):

that's right: you'll add up those 11 products. Then y ou'll multiply this sum by (delta x)/3, remembering that delta x is 0.2. OK?

OpenStudy (anonymous):

Yes. I did it and got the answer^_^

OpenStudy (mathmale):

Note that in using Simpson's Rule, n must be EVEN, and n+1 must be ODD. Our work here obeys those guidelines. Cool! think: How would this work be different if we were to use midpoints instead of subinterval endpoints as we just did with Simpson's Rule?

OpenStudy (mathmale):

One hint: We won't need coefficients such as 1, 4, 2, 4, ...., 4, 1.

OpenStudy (mathmale):

Lala?

OpenStudy (mathmale):

Into how many strips are we dividing the area under the graph of f(x) between 0 and 2?

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

My mistake. It's 12

OpenStudy (mathmale):

fOR Simpston's Rule it'd be 10 strips of width 0.2, and we'd need 10+1, or 11, y-values. but when using the Midpoint Rule, we need just 10 strips, each of width 0.2. We calculate the height of each strip and multiply each height by 0.2. Why?

OpenStudy (anonymous):

I thought the width is 0.1?

OpenStudy (mathmale):

No, the width is (2-0)/10, as before. When you typed in 0.1, you were referring to the midpoint of the first strip. The first strip begins at 0.0 and ends at 0.2, since the strip width is 0.2; the midpoint of the interval [0,0.2] is indeed 0.1.

OpenStudy (mathmale):

Please write out the 10 midpoints. Start with 0.1. For each new midpoint, simply add the width 0.2 to the previous midpoints. The 10 midpoints are then : ?

OpenStudy (anonymous):

0.1,0.3,0.5,0.7,0.9,1.1,1.3,1.5,1.7,1.9

OpenStudy (anonymous):

midpoint formula: (delta x/2)(f(0.1)+f(0.3)+......+f(1.9). Am I correct?

OpenStudy (mathmale):

Looks good. So, you'll be calculating 10 (not 11) function values, multiplying each by the strip width, and then adding up the 10 products. That's it. Less work than Simpson's Rule, but usually not as accurate as simpson's Rule. You OK with this? Any questions about the processes we've gone thru?

OpenStudy (anonymous):

Nope! Thank you so much^_^

OpenStudy (mathmale):

It's been a great pleasure to work with you, seeing that you already know a great deal about these topics! Bye, Lala. MM

OpenStudy (anonymous):

Bye Bye